# GSPRec: Temporal-Aware Graph Spectral Filtering for Recommendation

### v1

@@ 1 Introduction
Graph-based recommender systems are constrained by two core problems: using low-pass spectral filtering often suppresses distinct user signals in favor of broad popularity patterns, and the order of user interactions is typically neglected in their graph structures. Therefore, recommender systems should leverage both collaborative patterns and the rich temporal signals embedded within these interaction sequences [19]. Graph-based Collaborative Filtering (CF) methods [14, 42, 49] have achieved strong performance, but they face two limitations. First, models such as LightGCN [14], which rely on linear embedding propagation, act as low-pass filters [42, 33] by smoothing signals across the user-item graph, which encourages popularity but obscures user-specific patterns critical for personalization [3]. Second, while some GCN-based recommenders incorporate temporal data, they do so via separate sequential modeling components [26, 16] or by methods that do not integrate order into the graph structure for spectral analysis, which means they still discard sequential information needed to distinguish user preferences from popularity patterns [25, 42, 52, 38, 55].
@@
The filtering behavior inherent in many GCNs motivates a closer look through the lens of Graph Signal Processing (GSP), which generalizes classical signal processing to data defined on irregular graph structures [43, 34]. A core technique in GSP is spectral filtering, which transforms graph signals using the eigenbasis of the graph Laplacian [20]. Lowpass filters, which retain only low-frequency components, are widely adopted under the assumption that graph signals are typically smooth [43, 42, 14]. However, mid- and high-frequency components often encode task-relevant, localized, or personalized variations [7, 46, 17] motivating the design of spectral filters that can selectively amplify non-smooth components. Despite GSP’s potential, its adoption in recommender systems [25, 53, 52, 42] has largely focused on lowpass filtering to capture popular preference patterns, overlooking user-specific patterns.
@@
To overcome these limitations, we introduce GSPRec, a novel GSP-based framework. GSPRec achieves temporal awareness in spectral recommenders through two components: First, its sequentially-aware graph construction encodes multi-hop item transition patterns from user histories via diffusion into a symmetric graph structure, making temporal dynamics an intrinsic property for spectral analysis. Second, this graph empowers a frequency-aware dual-filtering, featuring a Gaussian bandpass filter to isolate vital mid-frequency user-specific preferences, complemented by a low-pass filter for popularity patterns.
@@
In this work, we encode the interaction graph with item-to-item transitions derived from user interaction sequences. A multi-hop diffusion process with exponential decay captures sequential dependencies while preserving the benefits of symmetric graph structures. These symmetric structures ensure real eigenvalues and orthogonal eigenvectors [6], which are crucial for stable spectral decomposition and well-defined filtering operations [34]. Unlike prior work that infers item-item similarity from collaborative information [53], our approach aggregates explicit sequential signals into the graph topology. Figure 1 illustrates this shift from sequence-based to graph-based modeling.
@@
Moreover, we design a Gaussian bandpass filter tailored to recommendation graphs. Informed by empirical spectral energy distributions [38, 46], the filter selectively captures mid-frequency components that are often associated with user-specific and sequential patterns positioned between low-frequency popularity signals and high-frequency noise. The Gaussian profile provides optimal localization in both spectral and vertex domains [43, 34], which enables smooth transitions across frequencies and reduces ringing artifacts commonly introduced by sharp filter cutoffs [12]. We summarize our contributions as follows:
@@
Figure 1: (a) Unidirectional transitions in models like RNNs [15, 18] process items in strict temporal order. (b) Bidirectional models such as BERT4Rec [45] allow context flow in both directions, but only between adjacent items. (c) Undirected item graph from user sequences. Solid lines show the initial symmetric graph $\mathbf{S}^{\prime}$. Orange dashed lines indicate higher-order relationships learned via multi-hop diffusion (Eq. 2) on $\mathbf{S}^{\prime}$, which aggregates weighted sequential paths ($\alpha$ and $d$). The strength of sequence-based connections between items, derived from the aggregated weighted paths of the diffusion process, is reflected in these symmetric links, making the graph ready for spectral filtering. This sequence is a running example to illustrate graph construction and transition modeling.
@@
2 Related Work & Background
@@
NGCF [49] General User-Item
LightGCN [14] General User-Item
LR-GCCF [4] General User-Item
IMP-GCN [23] General User-Item
@@
GF-CF [42] General LP User-Item
PGSP [25] General LP User-Item
HiGSP [53] General LP User-Item
FaGSP [52] General LP User-Item
Non-Propagative / Simplified Methods:
SimpleX [28] General User-Item
UltraGCN [29] General ApproxLP User-Item
@@
BERT4Rec [45] Sequential Sequence
SASRec [18] Sequential Sequence
@@
Graph-Based Collaborative Filtering:
Graph-based collaborative filtering represents users and items as nodes in a bipartite graph, modeling interactions through various approaches [14, 49, 55, 60, 17, 59]. Recent works address challenges including over-smoothing [3, 42], popularity bias [58], representation collapse [36], fairness [2, 56], and efficiency [22, 51], while others improve representations through disentanglement [50, 39, 9] or architectural simplification [29, 28]. Temporal dynamics have been incorporated through sequential models [18, 45, 21] or temporal graphs [57, 35]. Our approach encodes temporal locality directly as edge-level structure and applies spectral filtering without message passing.
Background on Graph Signal Processing:
Graph Signal Processing (GSP) analyzes signals on graph structures [34, 43, 40]. A graph signal maps values to nodes $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$. The normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ decomposes into eigenvalues $\boldsymbol{\Lambda}$ and eigenvectors $\mathbf{U}$ [6], forming the graph Fourier basis. Spectral filtering transforms signals:
$\mathbf{x}^{\prime}=\mathbf{U}\,g(\boldsymbol{\Lambda})\,\mathbf{U}^{T}\mathbf{x},\quad\text{where }g(\boldsymbol{\Lambda})=\mathrm{diag}(g(\lambda_{1}),\ldots,g(\lambda_{n}))$ (1)
Low-pass filters preserve global patterns, while band-pass filters capture mid-frequency variations related to user-specific behavior [34, 12]. In our context, user preference signals and item transition patterns enable spectral analysis as an alternative to neighborhood aggregation.
Graph Signal Processing in Recommendation:
Prior work has adapted GSP to recommender systems by treating interactions as graph signals [17, 27, 25, 52, 53, 61, 38]. These approaches employ low-pass filtering for global trends, neglecting mid-frequency components where user-specific patterns often reside. Various methods integrate spectral principles with neural architectures [52, 48, 54] or explore parameter-free formulations [24]. Our approach combines band-pass and low-pass filtering with temporal structure encoding, drawing from classical GSP filter design [20, 46, 30].
@@ 3 GSPRec
Figure 2: Overview of GSPRec. Starting from user-item interaction sequences (left), we construct a unified graph joining user-item links and item-item transitions via multi-hop diffusion with exponential decay (Eq. 2). This graph is processed using two complementary spectral filters: a band-pass filter to extract mid-frequency, user-specific patterns, and a low-pass filter to capture global trends.
GSPRec consists of a sequential graph encoding procedure and a frequency-aware spectral filtering framework. These components jointly capture two complementary signal patterns in the spectral domain: (1) globally consistent preference trends across the entire user-item graph, and (2) user-specific variation patterns that distinguish individual preferences. Figure 2 summarizes the pipeline from user-item interactions to final recommendations via graph construction and spectral filtering.
@@ 3.1 Problem Formulation
Let $\mathcal{U}=\{u_{1},\ldots,u_{m}\}$ denote users and =$\mathcal{I}=\{i_{1},\ldots,i_{n}\}$ denote items. Historical interactions are triplets $\mathcal{D}=\{(u,i,t)\mid u\in\mathcal{U},i\in\mathcal{I},t\in\mathbb{R}^{+}\}$, yielding binary matrix $\mathbf{X}\in\{0,1\}^{m\times n}$ where $x_{ui}=1$ if user $u$ interacted with item $i$. For each user $u$, we define a time-ordered sequence $\mathcal{S}_{u}=[i_{1},i_{2},\ldots]$ where $(u,i_{j},t_{j})\in\mathcal{D}$ and $t_{1}<t_{2}<\cdots$. Our task is to predict future user-item interactions by jointly modeling collaborative patterns and sequential dynamics. Specifically, we aim to predict unobserved entries in $\mathbf{X}$ using a unified spectral graph approach that leverages both user preference similarities and temporal item transitions.
@@ 3.2 Recommendation Graph Construction
Our goal is to construct a unified graph that integrates collaborative filtering signals with sequential transition patterns. This allows us to employ spectral methods to model both long-term user preferences and short-term sequential behavior, which typically require symmetric graph Laplacians [6]. The core challenge is to meaningfully translate inherently directed user sequences into a symmetric item-item graph component, denoted $\tilde{\mathbf{S}}$, suitable for this unified spectral analysis.
To achieve this, we shift from traditional sequence modeling paradigms to a graph-based representation that makes temporal dynamics amenable to spectral analysis. Figure 1 illustrates this conceptual shift: unlike models that process sequences directly through (a) unidirectional or (b) locally bidirectional mechanisms, our approach transforms user interaction sequences into (c) an undirected item graph. This graph is constructed to embed sequence-derived item proximity into a symmetric structure, a critical step for enabling spectral methods that typically require such symmetry. Although sequential models excel at capturing ordered context, our graph-based formulation is designed for compatibility with spectral analysis on symmetric matrices. This enables joint analysis of collaborative and sequential signals and decomposition with well-defined spectral properties [20]. However, using directed graphs directly with spectral methods would necessitate asymmetric Laplacians [5], which have different theoretical/computational aspects beyond our current scope. To create the symmetric item-item graph component $\tilde{\mathbf{S}}$ from sequences, we follow the following process.
@@
We define an initial directed transition matrix $\mathbf{S}\in\{0,1\}^{n\times n}$, where $s_{ij}=1$ if item $i$ directly precedes item $j$ in any user sequence. To adapt this for symmetric spectral methods, we form its undirected version, $\mathbf{S}^{\prime}\in\{0,1\}^{n\times n}$. In $\mathbf{S}^{\prime}$, an undirected edge $(i,j)$ exists if a transition $i\rightarrow j$ or $j\rightarrow i$ is in $\mathbf{S}$. This symmetrization is a standard practice when preparing directed graph for algorithms requiring undirected structures [47, 41]. While simplifying transition directionality, this approach allows capturing broader item relatedness from sequences, as rigid order is not always optimal and wider context can be beneficial [45]. Multi-hop diffusion (Eq. 2) operates on $\mathbf{S}^{\prime}$ to encode sequence-derived item associations.
@@
While $\mathbf{S}^{\prime}$ captures direct sequential links to model broader item influence propagated through multi-step paths within this symmetrized sequence-derived structure, we apply multi-hop diffusion directly to the symmetric graph $\mathbf{S}^{\prime}$:
@@
where $\alpha\in(0,1)$ is an exponential decay factor that gives higher weight to shorter paths in $\mathbf{S}^{\prime}$, thus emphasizing local sequential relationships. $d$ is the diffusion depth to control the extent of influence propagation. Since $\mathbf{S}^{\prime}$ is symmetric, each power $(\mathbf{S}^{\prime})^{k}$ is symmetric, and therefore the diffused matrix $\mathbf{S}^{(d)}$ is also symmetric. The entries $S^{(d)}_{ij}$ represent a score of multi-hop, locality-aware proximity between items $i$ and $j$, learned from the aggregated sequential patterns as illustrated in Figure 3. This diffusion process is analogous to using graph diffusion kernels to define similarity [10] and converges when $\alpha<1/\rho(\mathbf{S}^{\prime})$.
@@
The diffusion process as $d\to\infty$ is $\mathbf{S}^{(\infty)}=\sum_{k=1}^{\infty}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$. We can rewrite this as $\mathbf{S}^{(\infty)}=\mathbf{S}^{\prime}\sum_{j=0}^{\infty}(\alpha\mathbf{S}^{\prime})^{j}$ by letting $j=k-1$. This series converges if the standard matrix geometric series $\sum_{j=0}^{\infty}M^{j}$, where $M=\alpha\mathbf{S}^{\prime}$, converges. This occurs if the spectral radius $\rho(M)<1$, i.e., $\rho(\alpha\mathbf{S}^{\prime})<1$. Since $\rho(\alpha\mathbf{S}^{\prime})=\alpha\rho(\mathbf{S}^{\prime})$ (for $\alpha>0$), convergence requires $\alpha<1/\rho(\mathbf{S}^{\prime})$. ∎
Figure 3: Multi-hop diffusion (Eq. 2) on the symmetric base graph $\mathbf{S}^{\prime}$ (derived from initial directed sequences) with $\alpha=0.5$. Top: $\mathbf{S}^{\prime}$. Middle: Powers $(\mathbf{S}^{\prime})^{k}$. Bottom: $\mathbf{S}^{(3)}$ encoding decayed proximity.
In practice, we estimate $\rho(\mathbf{S}^{\prime})$ (e.g., via power iteration) and choose a suitable $\alpha$.
@@
This standard normalization technique [6] yields $\tilde{\mathbf{S}}$, ensuring its weights are scaled appropriately for stable spectral analysis. The weights in $\tilde{\mathbf{S}}$ reflect a transformation of user sequences into measures of item-item sequential relevance.
@@
We integrate the user-item interactions $\mathbf{X}$ and the sequence-derived item-item graph $\tilde{\mathbf{S}}$ into a single, symmetric adjacency matrix for the unified graph:
@@
This matrix $\mathbf{A}$ forms a heterogeneous graph where the $\mathbf{0}_{m\times m}$ block denotes no direct user-user links are explicitly modeled here, $\mathbf{X}$ captures user-item interactions, and $\tilde{\mathbf{S}}$ encodes the sequence-derived item-item relationships. From $\mathbf{A}$, we compute the symmetric normalized Laplacian, which is the cornerstone of our spectral filtering:
@@
The properties of this Laplacian are discussed in Proposition 2. Our approach, by constructing $\tilde{\mathbf{S}}$ from symmetrized sequential data and then applying diffusion, embeds sequence-derived proximity into a standard spectral framework, offering a distinct way to leverage temporal signals compared to pure sequential models that maintain strict causality [15, 18, 45].
@@
This follows from spectral graph theory [6], as $\mathbf{A}$ is constructed to be symmetric and non-negative, and Eq. 6 defines the standard normalized Laplacian. ∎

### v2

@@ 1 Introduction
Graph-based recommender systems are constrained by two core problems: using low-pass spectral filtering often suppresses distinct user signals in favor of broad popularity patterns, and the order of user interactions is typically neglected in their graph structures. Therefore, recommender systems should leverage both collaborative patterns and the rich temporal signals embedded within these interaction sequences Koren (2009). Graph-based Collaborative Filtering (CF) methods He et al. (2020); Shen et al. (2021); Wang et al. (2019) have achieved strong performance, but they face two limitations. First, models such as LightGCN He et al. (2020), which rely on linear embedding propagation, act as low-pass filters Shen et al. (2021) by smoothing signals across the user-item graph. This encourages broad popularity effects (e.g., everyone buys laptops) but suppresses subgroup-specific patterns (e.g., gamers buy laptops with mice, artists buy laptops with tabletsShen et al. (2021); Xia et al. (2025). Second, while some GCN-based recommenders incorporate temporal data, they do so via separate sequential modules Ma et al. (2020); Hsu & Li (2021) or through methods that do not integrate order into the graph structure for spectral analysis, which means they fail to preserve sequential information essential for distinguishing user preferences from popularity patterns.
@@
The filtering behavior inherent in many GCNs motivates a closer look through the lens of graph signal processing (GSP), which generalizes classical signal processing to irregular graphs Ortega et al. (2018). A core technique in GSP is spectral filtering, which transforms graph signals using the eigenbasis of the graph Laplacian Kruzick & Moura (2017). While low-pass filters, widely adopted under the assumption of smoothness Shen et al. (2021); He et al. (2020), capture global popularity, they fail to retain the more distinctive user-specific variations Liu et al. (2023); Xia et al. (2025). Recent methods have explored amplifying high frequencies Guo et al. (2023), but risk magnifying noise. By contrast, mid-frequency components encode localized, task-relevant variations Dong et al. (2019) remains underexplored, motivating filters that explicitly target this band for personalization.
@@
We introduce GSPRec, a novel GSP-based framework that achieves temporal awareness in spectral recommenders through two components: First, its sequentially-aware graph construction encodes multi-hop item transition patterns from user histories via diffusion into a symmetric graph structure to make temporal dynamics an intrinsic property for spectral analysis. Second, this graph empowers a frequency-aware dual-filtering by using Gaussian bandpass filter to extract mid-frequency user-specific preferences, complemented by a low-pass filter for popularity patterns.
@@
In this work, we encode the interaction graph with item-to-item transitions derived from user interaction sequences. A multi-hop diffusion process with exponential decay captures sequential dependencies while preserving the benefits of symmetric graph structures. The symmetric structure guarantees real eigenvalues and orthogonal eigenvectors Chung (1997), which are important for stable spectral decomposition and well-defined filtering operations Ortega et al. (2018). Unlike prior work that extracts item-item similarity from collaborative information Xia et al. (2024), our approach aggregates explicit sequential signals into the graph topology [^1] We note that sequences are used for graph construction, not for sequential next-item prediction.. Figure 1 illustrates this shift from sequence-based to graph-based modeling.
@@
Moreover, we design a Gaussian bandpass filter tailored to recommendation graphs. Informed by empirical spectral energy distributions Qin et al. (2024); Tremblay et al. (2018), the filter selectively captures mid-frequency components that are often associated with user-specific and sequential patterns positioned between low-frequency popularity signals and high-frequency noise. The Gaussian profile provides optimal localization in both spectral and vertex domains Shuman et al. (2013); Ortega et al. (2018), which enables smooth transitions across frequencies and reduces ringing artifacts commonly introduced by sharp filter cutoffs Hammond et al. (2011). We summarize our contributions as follows:
@@
Figure 1: Conceptual shift from sequence-based to graph-based modeling. (a) Unidirectional transitions in RNNs Hidasi et al. (2015); Kang & McAuley (2018). (b) Bidirectional models like BERT4Rec Sun et al. (2019). (c) Our undirected item graph construction: solid lines show initial symmetric graph $\mathbf{S}^{\prime}$, dashed lines show higher-order relationships from multi-hop diffusion.
@@
2 Background & Related Work
Graph Signal Processing Background:
Graph Signal Processing (GSP) analyzes signals on graph structures Ortega et al. (2018); Shuman et al. (2013). A graph signal maps values to nodes $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$. The normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ decomposes into eigenvalues $\bm{\Lambda}$ and eigenvectors $\mathbf{U}$ Chung (1997), enabling spectral filtering:
$\mathbf{x}^{\prime}=\mathbf{U}\,g(\bm{\Lambda})\,\mathbf{U}^{T}\mathbf{x},\quad\text{where }g(\bm{\Lambda})=\mathrm{diag}(g(\lambda_{1}),\ldots,g(\lambda_{n}))$ (1)
Low-pass filters preserve global patterns, while band-pass filters capture mid-frequency variations related to user-specific behavior Ortega et al. (2018); Hammond et al. (2011).
Graph-Based Collaborative Filtering:
Graph-based collaborative filtering represents users and items as nodes in a bipartite graph, modeling interactions through various approaches He et al. (2020); Wang et al. (2019); Ying et al. (2018); Zheng et al. (2018). Recent works address over-smoothing Chen et al. (2020a); Shen et al. (2021), popularity bias Zhang et al. (2023), and efficiency Liang et al. (2024), while others incorporate temporal dynamics through sequential models or temporal graphs Kang & McAuley (2018); Sun et al. (2019); Zhang et al. (2024); Ou et al. (2025). Our approach encodes temporal locality as edge-level structure and applies spectral filtering without message passing.
GSP in Recommendation:
Prior work adapts GSP to recommender systems by treating interactions as graph signals. Most methods adopt low-pass filtering Huang et al. (2017); Liu et al. (2023); Xia et al. (2025; 2024), while recent work Guo et al. (2023) explores high-pass filtering, which may recover details but also amplify noise Shuman et al. (2013); Ortega et al. (2018). These approaches operate on static user-item graphs without sequential dynamics. In this work, we integrate sequential transitions into the graph before spectral analysis.
@@ 2 Background & Related Work
NGCF Wang et al. (2019) General User-Item
LightGCN He et al. (2020) General User-Item
LR-GCCF Chen et al. (2020b) General User-Item
IMP-GCN Liu et al. (2021) General User-Item
@@
GF-CF Shen et al. (2021) General LP User-Item
JGCF Guo et al. (2023) General HP User-Item
PGSP Liu et al. (2023) General LP User-Item
HiGSP Xia et al. (2024) General LP User-Item
FaGSP Xia et al. (2025) General LP User-Item
@@
BERT4Rec Sun et al. (2019) Sequential Sequence
SASRec Kang & McAuley (2018) Sequential Sequence
@@
Table 1 summarizes existing approaches. Unlike GCN-based methods that rely on message passing, or GSP-based methods that apply low-pass (LP) or high-pass (HP) filtering to graphs, we join temporal graph construction (UI+Seq) with dual spectral filtering (BP+LP) to target mid-frequency components.
@@ 3 GSPRec
GSPRec consists of a sequential graph encoding procedure and a frequency-aware spectral filtering framework. These components jointly capture two complementary signal patterns in the spectral domain: (1) popularity patterns across the entire user-item graph, and (2) personalized patterns that distinguish individual preferences. Figure 2 summarizes the pipeline from user-item interactions to final recommendations via graph construction and spectral filtering.
Figure 2: Overview of GSPRec. Starting from user-item interaction sequences (left), we construct a unified graph joining user-item links and item-item transitions via multi-hop diffusion with exponential decay (Eq. 2). This graph is processed using two complementary spectral filters: a band-pass filter to extract mid-frequency, user-specific patterns, and a low-pass filter to capture popularity.
@@ 3.1 Problem Formulation
Let $\mathcal{U}=\{u_{1},\ldots,u_{m}\}$ denote users and $\mathcal{I}=\{i_{1},\ldots,i_{n}\}$ denote items. Historical interactions are triplets $\mathcal{D}=\{(u,i,t)\mid u\in\mathcal{U},i\in\mathcal{I},t\in\mathbb{R}^{+}\}$, yielding binary matrix $\mathbf{X}\in\{0,1\}^{m\times n}$ where $x_{ui}=1$ if user $u$ interacted with item $i$. For each user $u$, we define a time-ordered sequence $\mathcal{S}_{u}=[i_{1},i_{2},\ldots]$ where $(u,i_{j},t_{j})\in\mathcal{D}$ and $t_{1}<t_{2}<\cdots$. Our task is to predict future user-item interactions by jointly modeling collaborative patterns and sequential dynamics. We aim to predict unobserved entries in $\mathbf{X}$ using a unified spectral graph approach that leverages both user preference similarities and temporal item transitions.
@@ 3.2 Recommendation Graph Construction
Our goal is to construct a unified graph that integrates collaborative filtering signals with sequential transition patterns. Graph signal processing requires symmetric Laplacians to guarantee real eigenvalues, orthogonal eigenvectors, and stable spectral filtering Chung (1997); Ortega et al. (2018). Although user interaction histories are observed as timestamped sequences and thus appear directed, rigid order is not always reliable in recommendation domains Sun et al. (2019). We therefore transform sequences into a symmetric item-item graph component $\tilde{\mathbf{S}}$, a step that is theoretically required for spectral analysis. Multi-hop diffusion further preserves sequential proximity within this symmetric structure, allowing temporal information to remain embedded in the spectral representation.
Figure 1 illustrates our approach: while models in Figures 1(a) and 1(b) process sequences through uni/bidirectional mechanisms, we transform user interaction sequences into the undirected item graph in Figure 1(c). This enables joint analysis of collaborative and sequential signals with well-defined spectral properties Kruzick & Moura (2017). Direct use of asymmetric Laplacians Chung (2005) for directed graphs involves different theoretical formulation left for future work.
@@
We define an initial directed transition matrix $\mathbf{S}\in\{0,1\}^{n\times n}$, where $s_{ij}=1$ if item $i$ directly precedes item $j$ in any user sequence. To adapt this for symmetric spectral methods, we form its undirected version, $\mathbf{S}^{\prime}\in\{0,1\}^{n\times n}$. In $\mathbf{S}^{\prime}$, an undirected edge $(i,j)$ exists if a transition $i\rightarrow j$ or $j\rightarrow i$ is in $\mathbf{S}$. This symmetrization is a standard practice when preparing directed graph for algorithms requiring undirected structures Von Luxburg (2007); Satuluri & Parthasarathy (2011). While simplifying transition directionality, this approach allows capturing broader item relatedness from sequences, as rigid order is not always optimal and wider context can be beneficial Sun et al. (2019). Multi-hop diffusion (Eq. 2) operates on $\mathbf{S}^{\prime}$ to encode sequence-derived item associations.
[…]
@@
To model broader item influence through multi-step paths, we apply multi-hop diffusion to $\mathbf{S}^{\prime}$:
@@
where $\alpha\in(0,1)$ is an exponential decay factor that gives higher weight to shorter paths in $\mathbf{S}^{\prime}$, thus emphasizing local sequential relationships. $d$ is the diffusion depth to control the extent of influence propagation. Since $\mathbf{S}^{\prime}$ is symmetric, each power $(\mathbf{S}^{\prime})^{k}$ is symmetric, and therefore the diffused matrix $\mathbf{S}^{(d)}$ is also symmetric. The entries $S^{(d)}_{ij}$ represent a score of multi-hop, locality-aware proximity between items $i$ and $j$, learned from the aggregated sequential patterns as illustrated in Figure 3. This diffusion process is analogous to using graph diffusion kernels to define similarity Gasteiger et al. (2019) and converges when $\alpha<1/\rho(\mathbf{S}^{\prime})$. In practice, we estimate $\rho(\mathbf{S}^{\prime})$ (e.g., via power iteration) and choose a suitable $\alpha$.
@@
Figure 3: Multi-hop diffusion encoding item proximity from sequential patterns. Top: $\mathbf{S}^{\prime}$. Middle: Powers $(\mathbf{S}^{\prime})^{k}$. Bottom: $\mathbf{S}^{(3)}$ encoding decayed proximity.
@@
The diffusion process as $d\to\infty$ is $\mathbf{S}^{(\infty)}=\sum_{k=1}^{\infty}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$. We can rewrite this as $\mathbf{S}^{(\infty)}=\mathbf{S}^{\prime}\sum_{j=0}^{\infty}(\alpha\mathbf{S}^{\prime})^{j}$ by letting j=k1j=k-1. This series converges if the standard matrix geometric series $\sum_{j=0}^{\infty}M^{j}$, where $M=\alpha\mathbf{S}^{\prime}$, converges. This occurs if the spectral radius $\rho(M)<1$, i.e., $\rho(\alpha\mathbf{S}^{\prime})<1$. Since $\rho(\alpha\mathbf{S}^{\prime})=\alpha\rho(\mathbf{S}^{\prime})$ (for $\alpha>0$), convergence requires $\alpha<1/\rho(\mathbf{S}^{\prime})$. ∎
@@
This standard normalization technique Chung (1997) yields $\tilde{\mathbf{S}}$, ensuring its weights are scaled appropriately for stable spectral analysis. The weights in $\tilde{\mathbf{S}}$ reflect a transformation of user sequences into measures of item-item sequential relevance.
@@
We integrate the user-item interactions $\mathbf{X}$ and the sequence-derived item-item graph $\tilde{\mathbf{S}}$ into a single, symmetric adjacency matrix for the graph:
@@
This matrix $\mathbf{A}$ forms a heterogeneous unified graph where the $\mathbf{0}_{m\times m}$ block denotes no direct user-user links are explicitly modeled here, $\mathbf{X}$ captures user-item interactions, and $\tilde{\mathbf{S}}$ encodes the sequence-derived item-item relationships. From $\mathbf{A}$, we compute the symmetric normalized Laplacian, which is the cornerstone of our spectral filtering:
@@
The properties of this Laplacian are discussed in Proposition 2. Our approach, by constructing $\tilde{\mathbf{S}}$ from symmetrized sequential data and then applying diffusion, embeds sequence-derived proximity into a standard spectral framework, offering a distinct way to leverage temporal signals compared to pure sequential models that maintain strict causality Kang & McAuley (2018); Sun et al. (2019).
@@
This follows from spectral graph theory Chung (1997), as $\mathbf{A}$ is constructed to be symmetric and non-negative, and Eq. 6 defines the standard normalized Laplacian. ∎